# Measurement of reactor antineutrino oscillation at SNO+ M. Abreu Laboratório de Instrumentação e Física Experimental de Partículas (LIP), Av. Prof. Gama Pinto, 2, 1649-003, Lisboa, Portugal Universidade de Lisboa, Instituto Superior Técnico (IST), Departamento de Física, Av. Rovisco Pais, 1049-001 Lisboa, Portugal V. Albanese

### v1

@@ Introduction.
Several uncertainties on the flux are carried over from the detailed studies made by the Daya Bay experiment [20]. Per core, they include a 2.4% uncertainty from the isotopic emission spectra, 0.6% from fission fraction uncertainties, 0.5% from power output, and several other subdominant contributions. Studies carried out by the SNO+ collaboration produced consistent numbers.
@@
Antineutrinos up to about 3.3 MeV are produced by decays within the Earth. Only decays from the 238U and 232Th chains produce ’s with energy greater than the IBD threshold. Methods based on Ref. [25] are used to predict the expected geoneutrino flux at SNO+ in Terrestrial Neutrino Units (TNU), where 1 TNU corresponds to 1 IBD interaction per year per 1032 free protons. Assuming a constant survival probability of
@@
Across the dataset used in the present analysis, the model predicts
that 1.11.1 214BiPo coincidence events will pass the IBD selection criteria, which has a small effect on sensitivity estimates for [28].
@@
A study of simulated atmospheric neutrino interactions at SNO+ was performed using the GENIE Monte Carlo generator [29] and found a negligible contribution relative to the expected IBD signal.
@@
A binned extended likelihood fit was applied to the prompt energy distribution to find the best-fit values of and .
@@
The uncertainties on the prompt energy are dominated by reconstruction uncertainties. The energy systematics are modeled as uncertainties in the energy scale – both linear and non-linear – and in the energy resolution. Two independent factors parameterize the linear energy scale uncertainties of ’s/’s (1.8%) and protons (3%) to account for their different scintillation response. A nonlinear scaling following Birks’ law is applied as a variation in Birks’ constant (5.4%). Finally, energy resolution uncertainty is modeled as a Gaussian smearing with a standard deviation of 4.4% [28, 27].
A 2D visualization of the measurement significance is shown in Fig. 4 as a grid scan of the likelihood over and , profiling over all other parameters.
Performing a complete fit to account for correlations between all parameters, the best-fit values for the unconstrained oscillation parameters are = () 10-5 eV2 and = .
@@
Table 3 summarizes the results of the two different fits. All nuisance parameters are found to be in good agreement with expectations in both cases. Given that is largely determined by the flux and by the spectrum, the two parameters are minimally correlated, leaving little to gain by constraining one and fitting the other.
@@
Future sensitivities.
The sensitivity of this measurement is still limited by statistics and SNO+ is continuing to take data. The detector has been loaded with a secondary fluor (bis-MSB), increasing the light collected by more than 50%, and loading of tellurium is scheduled to begin in 2025. New data have been obtained with the enhanced light collection and sensitivity will remain similar when filled with tellurium, allowing continued study of reactor antineutrino oscillations and geoneutrinos.
Development of a novel event classifier showing improved rejection of the () background, as discussed in Ref. [7], is ongoing.
Because the () proton scatter events have an energy range similar to geo- IBDs, reduction of the () background is predicted to significantly improve the sensitivity to the geo- IBD rate.
@@
The SNO+ collaboration reports the second-most precise measurement of , using reactor antineutrinos. We find = () 10-5 eV2
while also measuring = . When applying Gaussian constraints according to external values of and , we obtain a measurement of = and of the geo- IBD rate: TNU.
@@ Acknowledgements.
For the purposes of open access, the authors have applied a Creative Commons Attribution licence to any Author Accepted Manuscript version arising. Representations of the data relevant to the conclusions drawn here are provided within this paper and its supplemental material.

### v2

@@ Introduction.
Several uncertainties on the flux are carried over from the detailed studies made by the Daya Bay experiment [20]. Per core, they include a 2.4% uncertainty from the isotopic emission spectra, 0.6% from fission fraction uncertainties, 0.5% from power output, and several other subdominant contributions. Studies carried out by the SNO+ collaboration produced consistent results.
@@
Antineutrinos up to about 3.3 MeV are produced by decays within the Earth. Only decays from the 238U and 232Th chains produce ’s with energy greater than the IBD threshold. Methods based on Ref. [25] are used to predict the geoneutrino flux at SNO+ (Sudbury, Ontario) in Terrestrial Neutrino Units (TNU), where 1 TNU corresponds to 1 IBD interaction per year per 1032 free protons. Assuming a constant survival probability of
@@
Across the dataset used in the present analysis, the model predicts that 1.11.1 214BiPo coincidence events will pass the IBD selection criteria, which has a small effect on sensitivity estimates for . This was also confirmed for different spectral shapes[28].
@@
A study of simulated atmospheric neutrino interactions at SNO+ was performed using the GENIE Monte Carlo generator
[29] and found a negligible contribution relative to the expected IBD signal.
@@
An extended likelihood fit to unbinned data was performed on the prompt energy distribution to find the best-fit values of and .
@@
The uncertainties on the prompt energy are dominated by reconstruction uncertainties [28, 27]. The energy systematics are modeled as uncertainties in the energy scale – both linear and non-linear – and in the energy resolution. Two independent factors parameterize the linear energy scale uncertainties of ’s/’s (1.8%) and protons (3%) to account for their different scintillation response and prediction uncertainties. A nonlinear scaling following Birks’ law is applied as a variation in Birks’ constant (5.4%). Finally, energy resolution uncertainty is modeled as a Gaussian smearing with a standard deviation of 4.4%.
A 2D visualization of the measurement significance is shown in Fig. 4 as a grid scan of the likelihood over and , profiling over all other parameters. Performing a complete fit to account for correlations between all parameters, the best-fit values for the unconstrained oscillation parameters are = () 10-5 eV2 and = .
@@
Table 2 shows the associated best-fit numbers of the signals and backgrounds.
@@
Table 3 summarizes the results of the two different fits. All nuisance parameters are found to be in good agreement with expectations in both cases.
Given that is largely determined by the flux and by the spectrum, the two parameters are minimally correlated, leaving little to gain by constraining one and fitting the other.
@@
Outlook.
Since the acquisition of the data used in this analysis, a secondary fluor (bis-MSB) was added to the SNO+ scintillator, increasing the detected light by more than 50%. The SNO+ search for neutrinoless double beta decay is scheduled to begin with the loading of tellurium near the end of 2025, and will allow the continued analysis of antineutrinos.
With its unique pattern of long-baseline reactor antineutrino oscillations, SNO+ will continue to update its independent measurement of , to compare and combine with those from dedicated experiments, such as KamLAND and JUNO [juno2022].
The SNO+ collaboration has identified a 214BiPo-like background component that is likely due to alpha-proton scattering and that has the potential to affect many experiments measuring IBDs. As more data are collected, this and the dominant () background will be better characterized. The use of pulse shape discrimination to suppress the () is in preparation and is predicted to significantly improve the experiment’s sensitivity to the geo- IBD rate.
With a larger flux of geoneutrinos expected from the thick North American plate, SNO+ measurements are highly complementary to observations at shallower locations, improving the separation of the crust and mantle fluxes most relevant for Earth modeling.
@@
With 286 tonne-years of data, the SNO+ collaboration reports the second-most precise measurement of , providing an independent evaluation of long-baseline reactor antineutrino oscillations. We find = () 10-5 eV2
while also measuring = .
Combining with previous measurements from long-baseline reactor antineutrinos and solar neutrinos yields = and a measurement of the geo- IBD rate at SNO+: TNU.